OpenStudy (anonymous):

Can someone help me with this please. Simplify the expression below. -7-3(t-5)=

OpenStudy (anonymous):

-7-3(t-5) -7-3t+15 8-3t=0 -3t=-8 then divide by -3 to isolate t t= 8/3

OpenStudy (anonymous):

-7-3t+15 -7+15=3t 8=3t t=8/3

OpenStudy (anonymous):

we beat him :P lol

OpenStudy (anonymous):

im still confuse

OpenStudy (anonymous):

distribute the -3 into (t-5)

OpenStudy (anonymous):

How would I write the whole equation out?

OpenStudy (anonymous):

then you'll get -7-3t+15=0

OpenStudy (anonymous):

now combine like terms which are -7 and 15 and that will give you 8

OpenStudy (anonymous):

I get it Now.